### v1

# Recursive Entropy in Thermodynamics: Establishing the Statistical-Physics Basis of the Zentropy Approach

Abstract
The recursive property of entropy is well known in the field of information theory; however, the concept is rarely used in the field of thermodynamics, despite being the field where the concept of entropy originated. This work shows that the equation for entropy used in the zentropy, which is an exact multiscale approach to thermodynamics, is a statement of the recursive property of entropy. Further, we clarify the meaning of entropy as the uncertainty arising from unconstrained degrees of freedom and separate configurational contributions from intra-configurational ones. Building on this, we derive the partition function, as used in zentropy, by maximizing entropy in its recursive form. The resulting framework is exact for a chosen level of description and enables principled coarse-graining, thereby reducing computational complexity while preserving thermodynamic consistency. These results position zentropy as a rigorous bridge between microscopic and macroscopic behavior, facilitating quantitative predictions and the study of emergent phenomena.
1 Introduction
In the present work, we set firm the mathematical foundations of the zentropy approach [Liu2024b, Liu2024c, Liu2022, Shang2023, Liu2019]. In discussions with colleagues, we have at times been met with bewilderment with regard to the entropy equation used in the zentropy approach:
$S=-k_{B}\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}.$ (1)
At first glance, there appears to be an extra term $\sum_{k}p_{k}S_{k}$ that does not belong since entropy is defined as
$S=-k_{B}\sum_{x}p_{x}\ln p_{x}.$ (2)
However, we will show that Eq. (1) is equivalent to Eq. (2) if we group outcomes $x$ into $k$ groups. Furthermore, there is a long history behind Eq. (1), and it has been readily used in the fields of information theory and quantum mechanics, where it is primarily known as the recursive property of entropy, which is a hierarchical form of the chain rule of entropy. The zentropy approach can be defined as the use of the recursive property to group entropy into more useful configurations, particularly in the context of quantum mechanics and density functional theory where zentropy was introduced.
The recursive property and chain rule are fundamental concepts in the fields of information theory and quantum information theory. They can be found at least as far back as 1938 [Tolman1938], and subsequently immortalized by Shannon in 1948 [Shannon1948]. However, the concept is rarely used in thermodynamics. It will be shown that restating entropy in the form of Eq. (1), enables a hierarchical coarse-graining approach to limit degrees of freedom, and greatly reduce computational complexity. The following textbooks are particularly helpful for general information on entropy and the recursive property. [Bengtsson2017, Nielsen2012, Cover2005]
In the remainder of the present work, we will (i) discuss the meaning of entropy and the recursive property, (ii) show an illustrative example of the recursive property of entropy to aid intuition, (iii) provide rigorous mathematical proof of the chain rule and the recursive property, (iv) derive the partition function based on recursive entropy as used in zentropy, and (v) discuss the usefulness of zentropy as an approach for emergent phenomena and reducing computational complexity. Note that in the following sections, we set the Boltzman constant equal to unity, $k_{B}=1$, so that temperature and energy share units. We anticipate that the present work will motivate members of the thermodynamics and condensed matter community to make use of the recursive property and zentropy.
2 The meaning of entropy and the recursive property
The combined first and second law of thermodynamics at equilibrium for closed systems under hydrostatic pressure may be written
dU=TdSPdVdU=TdS-PdV (3)
where temperature $T$ and pressure $P$ are potential quantities and entropy $S$ and volume $V$ are molar quantities. We intentionally avoid the use of intensive and extensive variables because normalized extensive variables are intensive [Hillert2007, Liu2024d, Liu2025] In many thermodynamics equations, $S$ and $V$ have symmetric places as molar quantities, and thermodynamics, per se, does not provide insight into the nature of volume and entropy. Instead, we turn to mechanics and statistical physics for answers. While volume is readily understood, entropy is more mysterious. Explaining entropy, first and foremost, we note that entropy is an anthropomorphic quantity [Jaynes1965]. It is a property of the measurement performed on the system. Entropy does not appear in Newton’s laws of motion, Maxwell’s equations, the Schrödinger equation, or Einstein’s field equations. Entropy only appears when there are degrees of freedom left unconstrained in the performance of a measurement. (While quantum mechanics contains measurement uncertainty due to wavefunction collapse, there is no uncertainty in the state of the wave function itself.) Specifically, entropy quantifies our uncertainty in the state of a system. It is important to note that this uncertainty is not from the inaccuracy of our instruments, but from unconstrained degrees of freedom of the system. When a measurement is performed on a system, usually, we do not fully constrain all degrees of freedom of the system. Instead, we reduce it to a statistical ensemble of possible configurations that comport with our measurement. For example, the canonical ensemble constrains the number of particles, volume, and temperature; it therefore includes all possible configurations compatible with these constraints regardless of other non-constrained degrees of freedom, say particle positions. All this is to explain that the entropy of the system depends on the constrained degrees of freedom. Now, what if we constrain a particular degree of freedom $k$ such that the entropy with the constraint is $S_{k}$ and the entropy without the constraint is $S$? The relation between the two is given exactly by Eq. (1).
3 Illustrative example
First, let us consider an illustrative example to intuitively understand what we mean by the recursive property using probability tree diagrams as shown in Fig. 1. [Bengtsson2017ch2]
Figure 1: Probability tree diagrams illustrating (a) the ungrouped and (b) grouped scenarios.
In the standard approach, we consider the probability vector $\vec{p}$ with components $p_{x}$, which describes the probability of each pure state. In this way, we write the entropy as,
$S(\vec{p})=-\sum_{x}p_{x}\ln p_{x}$ (4)
Now, instead of concerning ourselves with every possible outcome, let us group outcomes, and consider the probability distribution for each group using the probability vector $\vec{q}$ with components
$q_{k}=\sum_{x\in G_{k}}p_{x}$ (5)
where $G_{k}$ is the set of all the pure states for the $k$th group and each group is distinct so that $\{G_{k}\}$ forms a partition of the pure states (assumed here for intuitive ease, the proofs of Section 4 do not have this requirement). In this way, the entropy of the groups is
$S(\vec{q})=-\sum_{k}q_{k}\ln q_{k}$ (6)
This is often referred to as the configurational entropy in the literature [AdamGibbs1965, MiracleSenkov2017]. Of course, some entropy remains after choosing a particular group, and it is obvious that $S(\vec{q})\neq S(\vec{p})$. The entropy that remains after choosing a group is
$S_{k}\left(\frac{\vec{p}}{q_{k}}\right)=-\sum_{x\in G_{k}}\frac{p_{x}}{q_{k}}\ln\frac{p_{x}}{q_{k}}$ (7)
and represents the entropy that remains after constraining a particular degree of freedom $k$. We call this the intra-configurational entropy. The total entropy $S(\vec{p})$ can then be written as the configurational entropy and the weighted sum of the intra-configurational entropies,
$S(\vec{p})=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right)\\$ (8)
This can be shown using the fact that $p_{x}=q_{k}\cdot\frac{p_{x}}{q_{k}}$ and the logarithmic product rule $\ln{ab}=\ln a+\ln b$. Starting from the entropy of the pure states,
$S(\vec{p})$ $=-\sum_{x}p_{x}\ln p_{x}$ (9)
$=-\sum_{k}\sum_{x\in G_{k}}p_{x}\ln p_{x}$ (10)
$=-\sum_{k}\sum_{x\in G_{k}}p_{x}\ln\left(q_{k}\cdot\frac{p_{x}}{q_{k}}\right)$ (11)
$=-\sum_{k}\sum_{x\in G_{k}}p_{x}\left(\ln q_{k}+\ln\frac{p_{x}}{q_{k}}\right)$ (12)
$=-\sum_{k}\left(q_{k}\ln q_{k}+q_{k}\sum_{x\in G_{k}}\frac{p_{x}}{q_{k}}\ln\frac{p_{x}}{q_{k}}\right)$ (13)
$=-\sum_{k}\left(q_{k}\ln q_{k}-q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right)\right)$ (14)
$=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right)$ (15)
In this illustration, we partitioned pure-state configurations into groups. For this reason, the recursive property is sometimes called the grouping property. The justification for the recursive name is that we could just as easily group mixed-state configurations—group the groups, so to speak. Then the total entropy would be calculated recursively from subordinate scales.
4 Derivation of recursive entropy
To derive the entropy equation used in the zentropy approach, Eq. (1), we first use several definitions and theorems with the random variable formalization, as it facilitates conceptual progression. First, we start by defining the Shannon entropy [Cover2005]. Let $X$ be a random variable with outcomes $x$, alphabet $\mathcal{X}$, and probability mass function $p(x)$. For short, $p(x)=\Pr\{X=x\},x\in\mathcal{X}$.
Definition 1.
The Shannon entropy is defined as
$S(X)=-\sum_{x\in\mathcal{X}}p(x)\ln p(x)$ (16)
or equivalently
$S(p)=-\sum_{x}p_{x}\ln p_{x}$ (17)
where we define $0\ln 0\equiv 0$, a convention justified by noting that $\lim_{x\to 0}x\ln x=0$ and the intuition that a state $x$ with $p_{x}=0$ should not contribute to the entropy.
Next, we extend the definition of Shannon entropy to a pair of discrete random variables [Cover2005].
Definition 2.
The joint entropy of a pair of discrete random variables $(X,Y)$ with a joint distribution $p(x,y)$ is defined by
$S(X,Y)=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(x,y).$ (18)
Additionally, we consider the entropy of $Y$ conditional on knowing $X$. The entropy of $Y$ given $X$ is naturally the entropies of the conditional distributions averaged over the conditioning random variable [Cover2005].
Definition 3.
If the random pair $(X,Y)$ has a joint distribution $p(x,y)$ the conditional entropy is defined as
$S(Y\mid X)$ $=\sum_{x\in\mathcal{X}}p(x)S(Y\mid X=x)$ (19)
$=-\sum_{x\in\mathcal{X}}p(x)\sum_{y\in\mathcal{Y}}p(y\mid x)\ln p(y\mid x)$ (20)
$=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(y\mid x).$ (21)
We now present the first version of the chain rule, which states that the entropy of a pair of random variables can be expressed as the entropy of one variable plus the conditional entropy of the other [Cover2005].
Theorem 1 (Chain rule: two random variables).
$S(X,Y)=S(X)+S(Y\mid X)$ (22)
Proof.
Starting with Eq. (18) for the definition of joint entropy,
$S(X,Y)$ $=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(x,y)$ (23)
$=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln\left[p(x)p(y\mid x)\right]$ (24)
$=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(x)-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(y|x)$ (25)
$=-\sum_{x\in\mathcal{X}}p(x)\ln p(x)-\sum_{x\in\mathcal{X}}p(x)\sum_{y\in\mathcal{Y}}p(y\mid x)\ln p(y\mid x)$ (26)
$=S(X)+S(Y\mid X)\qed$ (27)
Following directly from the definitions of joint and conditional entropy, this proof illustrates how the chain rule naturally arises from these definitions. Because the definition of conditional entropy in Eq. (19) appears naturally in the final step, the chain rule formula is sometimes alternatively used as the definition of conditional entropy; see, for example, reference [Nielsen2012ch11].
For the second version of the chain rule, we generalize to $n$ random variables. By repeatedly applying the two-variable chain rule, the entropy of $n$ random variables can be expressed as the sum of conditional entropies by induction [Cover2005].
Theorem 2 (Chain rule: $n$ random variables).
$S(X_{1},X_{2},...,X_{n})=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1})$ (28)
Proof.
$S(X_{1},X_{2})$ $=S(X_{1})+S(X_{2}\mid X_{1})$ (29)
$S(X_{1},X_{2},X_{3})$ $=S(X_{1})+S(X_{2},X_{3}\mid X_{1})$ (30)
$=S(X_{1})+S(X_{2}\mid X_{1})+S(X_{3}\mid X_{2},X_{1})$ (31)
$=\vdots$ (32)
$S(X_{1},X_{2},...X_{n})$ $=S(X_{1})+S(X_{2}\mid X_{1})+...+S(X_{n}\mid X_{n-1},...X_{1})$ (33)
$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1})\qed$ (34)
Thus far, we have naturally treated each random variable on an equal level in accordance with the entropy being a symmetric function of the random variables. However, it is sometimes useful to consider one (or more) random variables hierarchically. That is, we wish to consider the entropy between configurations and the entropy within those configurations. To do this, we first define the entropy within a given configuration $k$:
Definition 4.
$S_{k}$ $=S(X_{1},X_{2},...X_{n}\mid K=k)$ (35)
$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1},K=k)$ (36)
Let $K$ be a random variable with probability distribution $p(K)$, where $\sum_{k}p_{k}=1$. We are now ready to prove our third version of the chain rule for the case of multiscale entropy.
Theorem 3 (Recursive property).
$S(K,X_{1},X_{2},...,X_{n})=-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}$ (37)
Proof.
First we choose to write the chain rule in the following form using the definition of conditional entropy from equation 19.
$S(X_{1},X_{2},...,X_{n})$ $=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1})$ (38)
$S(K,X_{1},X_{2},...,X_{n})$ $=S(K)+\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1},K)$ (39)
$=S(K)+\sum_{k\in\mathcal{K}}p(k)\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1},K=k)$ (40)
$=-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}\qed$ (41)
Note that in this form, there is a conspicuous absence of any reference to any of the $X_{i}$ on the right-hand side as they are all subsumed into $S_{k}$. Hence, the recursive property can be regarded as the hierarchical form of the chain rule.
The derivations presented in this section establish the mathematical foundation of the recursive property and show that the recursive property is not only consistent with the definition of entropy, but also essential. In fact, the recursive property is often taken as an axiom to single out the Shannon entropy as the natural choice of equations [khinchin1953concept, faddeev1956concept, Aczel1974] for entropy. The recursive property, as a hierachical form of the chain rule, reveal that the total entropy can be decomposed into a configuration component $\sum_{k}p_{k}\ln p_{k}$ and weighted sum of the intra-configurational entropies $\sum_{k}p_{k}S_{k}$. The recursive property represents a generalization of the standard entropy definition that reduces to the familiar form when configurations are chosen to be quantum pure states. That is, if $k$ specifies all degrees of freedom any other information ($X_{i}$) willbe redundant and the intra-configurational entropy is zero for all $k$, $\forall k\in\mathcal{K},\;S_{k}=0$. This flexibility in choosing the scale of description provides a powerful multiscale framework known as Zentropy for addressing systems on multiple scales where a natural partitioning of configurations exists, enabling efficient coarse-graining of degrees of freedom while maintaining thermodynamic consistency.
5 Derivation of the partition function
Definition 5.
The Helmholtz energy is defined as
F=ETSF=E-TS (42)
where the energy of the thermodynamic system is the expectation value of the fully specified configurations
$E=\langle E_{x}\rangle=\sum_{x}p_{x}E_{x}$ (43)
For the case where partially specified configurations form a partition, as we did in the illustrative example, by the law of total expectation, we know that $\langle E_{x}\rangle=\langle E_{k}\rangle$, and the energy of the system can also be expressed as the expectation value over the partially specified configurations.
$E=\sum_{k}p_{k}E_{k}$ (44)
As each configuration is now treated as an ensemble, we would expect $F_{k}$ to take the place of $E^{k}$ in the standard approach. Indeed, this is what we find. Substituting equations Eq. (44) and Eq. (37) into the definition of Helmholtz energy gives
$F=\sum_{k}p_{k}E_{k}+T\left(\sum_{k}p_{k}\ln p_{k}-\sum_{k}p_{k}S_{k}\right)$ (45)
Then by the definition of Helmholtz energy Def. 5, $F_{k}=E_{k}-TS_{k}$, and Eq. (45) simplies to
$F=\sum_{k}p_{k}F_{k}+T\sum_{k}p_{k}\ln p_{k}$ (46)
With the thermodynamic system energy and Helmholtz energies now given in terms of $E_{k}$ and $S_{k}$ we can now derive the partition function in terms of $E_{k}$ and $S_{k}$. Following the standard derivation of the partition function, we maximize entropy $S=-\sum_{k}p_{k}\ln p_{k}+\sum p_{k}S_{k}$ under the constraints that the sum of the probabilities is equal to one
$\sum_{k}p_{k}=1$ (47)
and the internal energy is a constant
$U=\langle E\rangle=\sum_{k}p_{k}E_{k}$ (48)
The Lagrange function is
$\mathcal{L}=\left(-\sum_{k}p_{k}\ln p_{k}+\sum p_{k}S_{k}\right)+\alpha\left(1-\sum_{k}p_{k}\right)+\beta\left(U-\sum_{k}p_{k}E_{k}\right)$ (49)
To extremize entropy we set the variation of the Lagrange function to zero
$0$ $=\delta\mathcal{L}$ (50)
$=\delta\left(-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}\right)+\delta\left(\alpha-\sum_{k}\alpha p_{k}\right)+\delta\left(\beta U-\sum_{k}\beta p_{k}E_{k}\right)$ (51)
$=\sum_{k}\left[\delta\left(-p_{k}\ln p_{k}+p_{k}S_{k}\right)-\delta\left(\alpha p_{k}\right)-\delta\left(\beta E_{i}p_{i}\right)\right]$ (52)
$=\sum_{k}\left[\frac{\partial}{\partial p_{k}}\left(-p_{k}\ln p_{k}\right)\delta p_{k}+\frac{\partial}{\partial p_{k}}\left(p_{k}S_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\alpha p_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\beta E_{i}p_{i}\right)\delta p_{k}\right]$ (53)
$=\sum_{k}\left[-\ln p_{k}-1+S_{k}-\alpha-\beta E_{i}\right]\delta p_{k}$ (54)
relying on the fact that entropy is symmetric with respect to $p_{k}$, for any variation
$0=-\ln p_{k}-1+S_{k}-\alpha-\beta E_{k}$ (55)
Isolating $p_{k}$ yields
$p_{k}=e^{\left(-1-\alpha-\beta E_{k}+S_{k}\right)}$ (56)
applying our probability constraint Eq.(47)
$1$ $=\sum_{k}p_{k}$ (57)
$=\sum_{k}e^{\left(-1-\alpha-\beta E_{k}+S_{k}\right)}$ (58)
$=\sum_{k}e^{-(1+\alpha)}e^{\left(-\beta E_{k}+S_{k}\right)}$ (59)
$=e^{-(1+\alpha)}\sum_{k}e^{\left(-\beta E_{k}+S_{k}\right)}$ (60)
$\frac{1}{e^{-(1+\alpha)}}=\sum_{k}e^{\left(-\beta E_{k}+S_{k}\right)}$ (61)
The partition function is then naturally defined as
$Z=\frac{1}{e^{-(1+\alpha)}}$ (62)
Substituting into Eq. (61) gives the relation between the partition function and $\beta$:
$Z=\sum_{k}e^{\left(-\beta E_{k}+S_{k}\right)}.$ (63)
While we have derived the partition function in terms of $E_{k}$ and $S_{k}$, in order to write it in terms of $F_{k}$, we assume here that $\beta$ is still inverse temperature then prove it subsequently.
$Z$ $=\sum_{k}e^{\left(-\frac{1}{T}E_{k}+S_{k}\right)}$ (64)
$=\sum_{k}e^{\left(-\frac{1}{T}(E_{k}-TS_{k})\right)}$ (65)
$=\sum_{k}e^{\left(-\frac{F_{k}}{T}\right)}=\sum_{k}e^{\left(-\beta F_{k}\right)}$ (66)
To confirm that the identity of $\beta$ is inverse temperature and is unaffected by the inclusion of intra-configurational entropy, we write $p_{k}$ in terms of the partition function in order to write entropy in terms of the partition function and finally apply the definition of temperature. Substituting Eq. (63) into eq. (56),
$p_{k}=\frac{1}{Z}e^{\left(-\beta E_{k}+S_{k}\right)}$ (68)
Substituting into the definition of the logarithmic portion of entropy
$S$ $=-\sum_{k}p_{k}\ln\left[\frac{1}{Z}e^{\left(-\beta E_{k}+S_{k}\right)}\right]+\sum_{k}p_{k}S_{k}$ (69)
$=-\sum_{k}p_{k}\left(-\beta E_{k}+S_{k}-\ln{Z}\right)+\sum_{k}p_{k}S_{k}$ (70)
$=-\sum_{k}p_{k}\left(-\beta E_{k}+S_{k}-\ln{Z}-S_{k}\right)$ (71)
$=-\sum_{k}p_{k}\left(-\beta E_{k}-\ln{Z}\right)$ (72)
$=-\sum_{k}\left(-p_{k}\beta E_{k}-p_{k}\ln{Z}\right)$ (73)
$=\beta\sum_{k}p_{k}E_{k}+\ln{Z}\sum_{k}p_{k}$ (74)
$=\beta E+\ln{Z}$ (75)
Differentiating with respect to energy gives
$\frac{\partial S}{\partial E}_{V,N}=\beta$ (77)
and by the definition of temperature $T=\frac{\partial E}{\partial S}_{V,N}$,
$\beta=\frac{1}{T}$ (78)
as expected.
Note that the definition of the partition function eq. (62) and the identity of $\beta$ are the same as in the standard thermodynamic approach, the derived relation between the partition function and $\beta$ is not. The difference being that $E_{x}$ is replaced by $F_{k}$, as expected.
6 Usefulness of zentropy
Why do we need recursive entropy when the definition is sufficient? Why group pure states? The answer is that recursive entropy is a generalization of the definition that provides a natural and efficiently computable approach to commonly encountered problems. Namely, problems where states of the system may be partitioned into basins. This partitioning into basins explains emergent pheonomena. Thus a framework, such as zentropy, that incorporates this partitioning of states, naturally predicts emergent phenomena.
The recursive property of entropy (Eq. 1) is a generalization of the definition. If the set of states are partitioned so finely that each $k$ already specifies a pure state (the number of $k$ is maximal), then $S_{k}$ is equal to zero for all $k$, only the first term $k_{B}\sum_{k}p_{k}\ln p_{k}$ remains, $k=x$, and we recover the definition of entropy (Eq. 2). Keeping the intra-configurational term therefore just allows us to choose configurations that contain additional unresolved mixed-up-ness. Recursive entropy is thus a generalization and reduces to the definition when no hidden uncertainty remains in each configuration ($S_{k}=0)$.
Conversely, at the opposite extreme, we could choose our set of states to be as coarsely partitioned as possible so that there is only one configuration (The number of $k$ is minimal), in which case, the configurational entropy term is zero $k_{B}\sum_{k}p_{k}\ln p_{k}=0$ and Eq. 1 becomes trivial $S=\sum_{k}p_{k}S_{k}=S_{k}$. However, it is often best to choose a partitioning that is somewhere inbetween. All this is to say that by including the intra-configurational entropy term $\sum_{k}p_{k}S_{k}$—using the recursive property of entropy— allows us to choose a definition of configuration that is most useful. Or in other words, to define the configurations of the system at the most useful scale.
A useful way to group pure states is in the same way as they are partitioned in nature, that is, to partition pure states in accordance with the basins of the energy landscape. This way, it is natural to consider the system’s behavior given a specifed basin/configuration. This can be used to simplify calculations by applying approximations hierarchically, as the pure states in each basin often have some shared properties. For example, consider a magnetic crystal, where each configuration is a given magnetic ordering. The remaining uncertainty is then in the phonon and electron states given that magnetic ordering.
$S_{k}=S_{k,\text{el}}+S_{k,\text{vib}}$ (79)
where $S_{k}$ is the entropy of a given configuration/ magnetic ordering, $S_{\text{el}}$ is entropy due to electron excitations, and $S_{\text{vib}}$ due to lattice vibrations. Practically, this may be calculated using density functional theory. The additivity of $S_{\text{el}}$ and $S_{\text{vib}}$ assumes that the phonon density of states and the electron density of states are independent for each given configuration. This assumption facilitates the calculations. Otherwise, they are subadditive, and an additional relation between $S_{\text{el}}$ and $S_{\text{vib}}$ is required to determine the $S_{k}$ exactly. However, such a calculation is difficult and often unnecessary.
Here we clarify some terminology. There are two different uses of the term ”coarse-graining”. To clarify, one meaning refers to the partitioning or, more generally, grouping of states, and the other refers to the incomplete sampling of configurations. While the term ”coarse-graing” is often used for both, The two meanings are very different. Notably, incomplete sampling is approximate, whereas grouping of states and calculating entropy with the recursive property is exact.
We now turn our attention to the incomplete sampling of configurations. While approximate, it is often not feasible to calculate all possible configurations of a system as required by the partition function. Instead, a representative sample of configurations may be used. For example, this may be achieved by calculating all the configurations of a finite cell rather than the infinite lattice, or using Monte Carlo or other methods to sample only configurations of sufficient probability. By grouping pure states, we can more efficiently sample or approximate the pure states. For example, by partitioning pure states by spin configuration, we can apply the harmonic approximation to each configuration to gather the phonon density of states.
Including both configurational and intra-configurational terms in the equation for entropy, allows defining configurations on a scale between the macroscopic thermodynamic state and quantum pure states. As mentioned earlier, for partitions, when the number of configurations $k$ is maximal, the configurations are defined on the scale of quantum pure states, and when the number of configurations is minimal, the configuration is the macroscopic thermodynamic state. By choosing the partition such that the number of configurations is between the two extremes, we build a bridge between scales.
This bridge can be useful in the description of emergent phenomena. While there are rules that govern the quantum pure states (quantum mechanics), one could form new rules that govern the configurations formed by the grouping. Thus, a rule set for the quantum pure states, and an emergent rule set for the configurations govern the same system. This approach works particularly well for configurations wherein the quantum pure states (or more gennerally subconfigurations) are strongly lumpable [Rosas2024]. That is, the quantum pure states of one group have a defining property that is clearly distinct from quantum pure states of another configuration.
In summary, the recursive property of entropy provides a flexible and exact framework for partitioning thermodynamic systems at physically meaningful scales. By grouping pure states according to natural basins in the energy landscape, we can apply hierarchical approximations that exploit shared properties within each configuration, simplifying otherwise intractable calculations. The multiscale nature of the approach naturally bridges microscopic and macroscopic descriptions. This bridging capability makes zentropy particularly powerful for describing emergent phenomena, where different rule sets govern behavior at different scales, provided the configurations are sufficiently distinct or ”strongly lumpable.” The result is a computationally practical approach that maintains rigorous statistical-mechanical foundations while enabling quantitative predictions across scales. Here, an analogy can be drawn between zentropy and Kohn–Sham DFT, which provides a practical approach to quantum mechanics by replacing an intractable interacting many-electron problem with a non-interacting system in an effective potential that preserves the exact density.
7 Conclusion
We have established a clear statistical-mechanical foundation for the zentropy approach by proving the recursive property of entropy as a hierarchical form of the chain rule and showing its equivalence to the standard definition when configurations reduce to pure states. Building on this foundation, we derived a grouped form of the Helmholtz energy and obtained a partition function expressed in terms of configuration Helmholtz energies. This zentropy approach separates configurational entropy from intra-configurational contributions and provides a principled, exact route to multiscale modeling, by aligning configurations with natural basins in the energy landscape
Practically, these results position zentropy as a computationally efficient bridge between microscopic states and macroscopic behavior, clarifying when and how emergent phenomena can be captured by a reduced description. We anticipate that adopting the recursive property in thermodynamics and condensed-matter calculations will streamline predictive workflows, particularly in DFT and when configurations are strongly lumpable.
8 Acknowledgments
The present work is funded by the Department of Energy (DOE) through Grant No. DE-SC0023185

### v2

@@
 The recursive property of entropy is well known in the field of information theory; however, the concept is rarely used in the field of thermodynamics, despite being the field where the concept of entropy originated.
-This work shows that the equation for entropy used in the zentropy, which is an exact multiscale approach to thermodynamics, is a statement of the recursive property of entropy.
-Further, we clarify the meaning of entropy as the uncertainty arising from unconstrained degrees of freedom and separate configurational contributions from intra-configurational ones.
-Building on this, we derive the partition function, as used in zentropy, by maximizing entropy in its recursive form.
+This work shows that the equation for entropy used in the zentropy approach, which is an exact multiscale approach to thermodynamics, is a statement of the recursive property of entropy.
+Further, we clarify the meaning of entropy as the uncertainty arising from unconstrained degrees of freedom and separate configurational and intra-configurational contributions.
+Building on this, we derive the Helmholtz energy and the partition function, as used in zentropy.
 The resulting framework is exact for a chosen level of description and enables principled coarse-graining, thereby reducing computational complexity while preserving thermodynamic consistency.
@@
 1 Introduction
-In the present work, we set firm the mathematical foundations of the zentropy approach [Liu2024b, Liu2024c, Liu2022, Shang2023, Liu2019].
-In discussions with colleagues, we have at times been met with bewilderment with regard to the entropy equation used in the zentropy approach:
+The Gibbs (Shannon) entropy is a function of the probabilities of all possible microstates of a system and quantifies the uncertainty in the state of the system,
+$S=-k_{B}\sum_{x}p_{x}\ln p_{x}.$ (1)
+From this definition, there is no prescription for the ordering of the microstates.
+The only requirement is that the summation be performed over the entire set.
+This leaves us with considerable freedom in our description of the system.
+In fact, as is done in the zentropy approach, the summation may even be performed hierarchically, with some adjustments.
+The zentropy approach [Liu2019, Shang2023, Hew2024, Hong2025] exploits this choice in description to facilitate hierarchical coarse-graining, and the approach was first introduced as a method of “multi-scale entropy” [Liu2019].
+Later, the method adopted the name zentropy [Liu2022], where “Z” refers to the partition function $Z$, itself derived from the German word Zustandssumme, meaning “sum over states”.
+The term zentropy emphasizes that this approach sums over states with entropy.
+[…]
+In the present work, we set firm and elaborate on the mathematical foundations of the zentropy approach, where entropy is given by
 $S=-k_{B}\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}.$ (2)
-At first glance, there appears to be an extra term $\sum_{k}p_{k}S_{k}$ that does not belong since entropy is defined as
-$S=-k_{B}\sum_{x}p_{x}\ln p_{x}.$ (2)
+At first glance, there appears to be an extra term $\sum_{k}p_{k}S_{k}$.
 However, we will show that Eq.
 (2) is equivalent to Eq.
-(2) if we group outcomes $x$ into $k$ groups.
+(1) if we group outcomes $x$ into groups $k$.
 Furthermore, there is a long history behind Eq.
-(1), and it has been readily used in the fields of information theory and quantum mechanics, where it is primarily known as the recursive property of entropy, which is a hierarchical form of the chain rule of entropy.
-The zentropy approach can be defined as the use of the recursive property to group entropy into more useful configurations, particularly in the context of quantum mechanics and density functional theory where zentropy was introduced.
+(2), where it is primarily known as the recursive property of entropy, which is a hierarchical form of the chain rule of entropy.
+This property of entropy has been widely used in the fields of information theory and quantum mechanics.
+The zentropy approach can be defined as the use of the recursive property to group microstates into more useful coarse-grained configurations, particularly in the context of quantum mechanics and density functional theory, where zentropy was introduced.
 The recursive property and chain rule are fundamental concepts in the fields of information theory and quantum information theory.
-They can be found at least as far back as 1938 [Tolman1938], and subsequently immortalized by Shannon in 1948 [Shannon1948].
+They can be found at least as far back as 1938 [Tolman1938, Chap.
+IV, p.
+176], and subsequently immortalized by Shannon in 1948 [Shannon1948].
 However, the concept is rarely used in thermodynamics.
 It will be shown that restating entropy in the form of Eq.
-(1), enables a hierarchical coarse-graining approach to limit degrees of freedom, and greatly reduce computational complexity.
-The following textbooks are particularly helpful for general information on entropy and the recursive property.
-[Bengtsson2017, Nielsen2012, Cover2005]
+(2), enables a hierarchical coarse-graining approach to limit degrees of freedom and greatly reduce computational complexity.
+The following textbooks are particularly helpful for general information on entropy and the recursive property [Bengtsson2017, Nielsen2012, Cover2005].
 In the remainder of the present work, we will (i) discuss the meaning of entropy and the recursive property, (ii) show an illustrative example of the recursive property of entropy to aid intuition, (iii) provide rigorous mathematical proof of the chain rule and the recursive property, (iv) derive the partition function based on recursive entropy as used in zentropy, and (v) discuss the usefulness of zentropy as an approach for emergent phenomena and reducing computational complexity.
-Note that in the following sections, we set the Boltzman constant equal to unity, $k_{B}=1$, so that temperature and energy share units.
-We anticipate that the present work will motivate members of the thermodynamics and condensed matter community to make use of the recursive property and zentropy.
-2 The meaning of entropy and the recursive property
-The combined first and second law of thermodynamics at equilibrium for closed systems under hydrostatic pressure may be written
-dU=TdSPdVdU=TdS-PdV (3)
-where temperature $T$ and pressure $P$ are potential quantities and entropy $S$ and volume $V$ are molar quantities.
-We intentionally avoid the use of intensive and extensive variables because normalized extensive variables are intensive [Hillert2007, Liu2024d, Liu2025] In many thermodynamics equations, $S$ and $V$ have symmetric places as molar quantities, and thermodynamics, per se, does not provide insight into the nature of volume and entropy.
-Instead, we turn to mechanics and statistical physics for answers.
+Note that in the following sections, we set the Boltzmann constant equal to unity, $k_{B}=1$, so that temperature and energy share the same units.
+We anticipate that the present work will motivate members of the thermodynamics and condensed matter communities to make use of the recursive property and zentropy.
+2 The Meaning of Entropy: Constrained and Unconstrained Degrees of Freedom
+For a simple compressible closed system at thermodynamic equilibrium, the fundamental differential relation may be written as
+dU=TdSPdV.dU=TdS-PdV. (3)
+In this form, $S$ and $V$ appear symmetrically as conjugate variables to $T$ and $P$, respectively, and thermodynamics per se does not provide insight into the nature of volume and entropy.
+Instead, we must rely on mechanics and statistical physics for answers.
 While volume is readily understood, entropy is more mysterious.
-Explaining entropy, first and foremost, we note that entropy is an anthropomorphic quantity [Jaynes1965].
+To explain entropy, first and foremost, we note that entropy is an anthropomorphic quantity [Jaynes1965].
 It is a property of the measurement performed on the system.
 Entropy does not appear in Newton’s laws of motion, Maxwell’s equations, the Schrödinger equation, or Einstein’s field equations.
-Entropy only appears when there are degrees of freedom left unconstrained in the performance of a measurement.
+Instead, it only appears when there are degrees of freedom left unconstrained when performing a measurement.
 (While quantum mechanics contains measurement uncertainty due to wavefunction collapse, there is no uncertainty in the state of the wavefunction itself.) Specifically, entropy quantifies our uncertainty in the state of a system.
-It is important to note that this uncertainty is not from the inaccuracy of our instruments, but from unconstrained degrees of freedom of the system.
-When a measurement is performed on a system, usually, we do not fully constrain all degrees of freedom of the system.
-Instead, we reduce it to a statistical ensemble of possible configurations that comport with our measurement.
-For example, the canonical ensemble constrains the number of particles, volume, and temperature; it therefore includes all possible configurations compatible with these constraints regardless of other non-constrained degrees of freedom, say particle positions.
-All this is to explain that the entropy of the system depends on the constrained degrees of freedom.
+It is important to note that this uncertainty is not due to instrument inaccuracy but to the unconstrained degrees of freedom of the system.
+When a measurement is performed on a system, we usually do not fully constrain all degrees of freedom of the system.
+Instead, we reduce it to a statistical ensemble of possible microstates that comport with our measurement.
+For example, the canonical ensemble constrains the number of particles, volume, and temperature; it therefore includes all possible microstates compatible with these constraints, regardless of other unconstrained degrees of freedom, such as particle positions.
+For clarity, “degrees of freedom” here refers to the microscopic degrees of freedom, not to the macroscopic state variable degrees of freedom of Gibbs’ phase rule.
+All this is to explain that the entropy of the system also depends on the constrained degrees of freedom.
 Now, what if we constrain a particular degree of freedom $k$ such that the entropy with the constraint is $S_{k}$ and the entropy without the constraint is $S$?
@@
 (2).
-3 Illustrative example
-First, let us consider an illustrative example to intuitively understand what we mean by the recursive property using probability tree diagrams as shown in Fig. 1. [Bengtsson2017ch2]
+3 Intuitive Illustration of Recursive Entropy: Partitioned Case
+To intuitively understand what we mean by the recursive property, let us consider an illustrative example using probability tree diagrams [Bengtsson2017ch2] as shown in Figure 1.
 Figure 1: Probability tree diagrams illustrating (a) the ungrouped and (b) grouped scenarios.
-In the standard approach, we consider the probability vector $\vec{p}$ with components $p_{x}$, which describes the probability of each pure state. In this way, we write the entropy as,
-$S(\vec{p})=-\sum_{x}p_{x}\ln p_{x}$ (4)
-Now, instead of concerning ourselves with every possible outcome, let us group outcomes, and consider the probability distribution for each group using the probability vector $\vec{q}$ with components
-$q_{k}=\sum_{x\in G_{k}}p_{x}$ (5)
-where $G_{k}$ is the set of all the pure states for the $k$th group and each group is distinct so that $\{G_{k}\}$ forms a partition of the pure states (assumed here for intuitive ease, the proofs of Section 4 do not have this requirement).
-In this way, the entropy of the groups is
-$S(\vec{q})=-\sum_{k}q_{k}\ln q_{k}$ (6)
-This is often referred to as the configurational entropy in the literature [AdamGibbs1965, MiracleSenkov2017].
+In the ungrouped scenario (Figure 1a), we consider the probability vector $\vec{p}$ with components $p_{x}$, which describes the probability of each microstate. In this way, we write the entropy as,
+$S(\vec{p})=-\sum_{x}p_{x}\ln p_{x}.$ (4)
+Now, instead of concerning ourselves with every possible outcome, let us group outcomes (Figure 1b), and consider the probability distribution for each group or partially-constrained configuration using the probability vector $\vec{q}$ with components $q_{k}$ defined by
+$q_{k}=\sum_{x\in G_{k}}p_{x},$ (5)
+where $G_{k}$ is the set of all fully-constrained configurations for the $k$th group and each group is distinct so that $\{G_{k}\}$ forms a partition of the fully-constrained configurations (assumed here for intuition; the proofs in Section 4 do not require strict partitions and hold in the more general case).
+In this way, the entropy between the groups is
+$S(\vec{q})=-\sum_{k}q_{k}\ln q_{k}.$ (6)
+This is often referred to as the configurational entropy in the literature [AdamGibbs1965, MiracleSenkov2017, Hong2025, Shang2023].
 Of course, some entropy remains after choosing a particular group, and it is obvious that $S(\vec{q})\neq S(\vec{p})$.
 The entropy that remains after choosing a group is
-$S_{k}\left(\frac{\vec{p}}{q_{k}}\right)=-\sum_{x\in G_{k}}\frac{p_{x}}{q_{k}}\ln\frac{p_{x}}{q_{k}}$ (7)
-and represents the entropy that remains after constraining a particular degree of freedom $k$.
+$S_{k}\left(\vec{p}_{|k}\right)=-\sum_{x\in G_{k}}\frac{p_{x}}{q_{k}}\ln\frac{p_{x}}{q_{k}},$ (7)
+where $\vec{p}_{|k}$ is the probability distribution after choosing a particular group $k$ and has components $p_{x|k}=p_{x}/q_{k}$.
+$S_{k}$ represents the entropy that remains after constraining a particular degrees of freedom.
 We call this the intra-configurational entropy.
-The total entropy $S(\vec{p})$ can then be written as the configurational entropy and the weighted sum of the intra-configurational entropies,
-$S(\vec{p})=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right)\\$ (8)
-This can be shown using the fact that $p_{x}=q_{k}\cdot\frac{p_{x}}{q_{k}}$ and the logarithmic product rule $\ln{ab}=\ln a+\ln b$. Starting from the entropy of the pure states,
+The total entropy $S(\vec{p})$ can then be written as the configurational entropy plus the weighted sum of the intra-configurational entropies
+$S(\vec{p})=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right).$ (8)
+This can be shown using the fact that $p_{x}=q_{k}\cdot\frac{p_{x}}{q_{k}}$ and the logarithmic product rule $\ln{ab}=\ln a+\ln b$. Starting from the entropy of the fully-constrained configurations
 $S(\vec{p})$ $=-\sum_{x}p_{x}\ln p_{x}$ (9)
@@
 $=-\sum_{k}\left(q_{k}\ln q_{k}-q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right)\right)$ (14)
-$=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right)$ (15)
-In this illustration, we partitioned pure-state configurations into groups.
+$=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right).$ (15)
+In this illustration, we partitioned fully-constrained configurations into groups.
 For this reason, the recursive property is sometimes called the grouping property.
-The justification for the recursive name is that we could just as easily group mixed-state configurations—group the groups, so to speak.
-Then the total entropy would be calculated recursively from subordinate scales.
-4 Derivation of recursive entropy
-To derive the entropy equation used in the zentropy approach, Eq.
-(1), we first use several definitions and theorems with the random variable formalization, as it facilitates conceptual progression.
-First, we start by defining the Shannon entropy [Cover2005].
+The term recursive is justified because we could just as easily group the partially-constrained configurations—group the groups, so to speak.
+Some terminological clarification is required here.
+In statistical mechanics, the terms microstate and macrostate are introduced to clarify the meaning of state.
+In thermodynamics, the state of a system has a property called entropy.
+However, the fully specified microscopic description of the same system is purely mechanical, and there are no states with entropy.
+Thus, the macrostate refers to the system as it appears macroscopically, with many degrees of freedom left unconstrained, while the microstate refers to the system fully specified, with all degrees of freedom constrained.
+This distinction becomes more subtle in the zentropy approach, in which the system has partially-constrained configurations.
+In this context, configurations may be viewed two ways: they can be microstates that comprise a macrostate or super-configuration, or they can be macrostates with their own microstates or sub-configurations.
+To avoid ambiguity, we use the term configuration to refer generally to anything that corresponds to a particular group $k$ of Eq.
+(2) (or Eq.
+(8) of this section).
+Further, a configuration may be called a microstate when treated as a sub-configuration and a macrostate when treated as a super-configuration.
+4 Derivation of Recursive Entropy: General Case
+While the previous section provided the intuition and showed the particular case where the fully-constrained configurations are partitioned, in this section, we provide a general proof of the recursive property that does not rely on forming a partition.
+To do this, we first establish several definitions and theorems, largely following the approach and notation of Cover and Thomas [Cover2005], starting with the definition of Shannon entropy.
 Let $X$ be a random variable with outcomes $x$, alphabet $\mathcal{X}$, and probability mass function $p(x)$.
@@
 The Shannon entropy is defined as
-$S(X)=-\sum_{x\in\mathcal{X}}p(x)\ln p(x)$ (16)
-or equivalently
-$S(p)=-\sum_{x}p_{x}\ln p_{x}$ (17)
-where we define $0\ln 0\equiv 0$, a convention justified by noting that $\lim_{x\to 0}x\ln x=0$ and the intuition that a state $x$ with $p_{x}=0$ should not contribute to the entropy.
-Next, we extend the definition of Shannon entropy to a pair of discrete random variables [Cover2005].
+$S(X)=-\sum_{x\in\mathcal{X}}p(x)\ln p(x),$ (16)
+where we define $0\ln 0\equiv 0$, a convention justified by noting that $\lim_{x\to 0}x\ln x=0$ and that the intuition that a microstate $x$ with $p(x)=0$ should not contribute to the entropy.
+Next, we extend the definition of Shannon entropy to a pair of discrete random variables.
+Let $Y$ also be a random variable with alphabet $\mathcal{Y}$ and probability mass function $p(y)=\Pr\{Y=y\}$, and let $p(x,y)$ denote the joint probability mass function of $(X,Y)$.
 Definition 2.
-The joint entropy of a pair of discrete random variables $(X,Y)$ with a joint distribution $p(x,y)$ is defined by
+The joint entropy of a pair of discrete random variables $(X,Y)$ with a joint probability mass function $p(x,y)$ is defined by
 $S(X,Y)=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(x,y).$ (17)
-Additionally, we consider the entropy of $Y$ conditional on knowing $X$.
-The entropy of $Y$ given $X$ is naturally the entropies of the conditional distributions averaged over the conditioning random variable [Cover2005].
+Additionally, we consider the entropy of $Y$ conditional on knowing $X$. This entropy is naturally the entropies of the conditional distributions averaged over the conditioning random variable.
 Definition 3.
-If the random pair $(X,Y)$ has a joint distribution $p(x,y)$ the conditional entropy is defined as
+If the random pair $(X,Y)$ has a joint probability mass function $p(x,y)$ the conditional entropy is defined as
 $S(Y\mid X)$ $=\sum_{x\in\mathcal{X}}p(x)S(Y\mid X=x)$ (18)
 $=-\sum_{x\in\mathcal{X}}p(x)\sum_{y\in\mathcal{Y}}p(y\mid x)\ln p(y\mid x)$ (19)
-$=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(y\mid x).$ (21)
-We now present the first version of the chain rule, which states that the entropy of a pair of random variables can be expressed as the entropy of one variable plus the conditional entropy of the other [Cover2005].
+$=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(y\mid x),$ (20)
+where the last equality follows from the definition of the joint probability mass function, $p(x,y)=p(x)p(y\mid x)$.
+We now present the first version of the chain rule, which states that the entropy of a pair of random variables can be expressed as the entropy of one variable plus the conditional entropy of the other.
 Theorem 1 (Chain rule: two random variables).
@@
 Proof.
-Starting with Eq. (18) for the definition of joint entropy,
+Starting from the definition of joint entropy (Eq. (17)),
 $S(X,Y)$ $=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(x,y)$ (22)
@@
 $=-\sum_{x\in\mathcal{X}}p(x)\ln p(x)-\sum_{x\in\mathcal{X}}p(x)\sum_{y\in\mathcal{Y}}p(y\mid x)\ln p(y\mid x)$ (25)
-$=S(X)+S(Y\mid X)\qed$ (27)
+$=S(X)+S(Y\mid X).\qed$ (26)
+In the fourth line of the derivation, we used the marginal probability relation $p(x)=\sum_{y\in\mathcal{Y}}p(x,y)$.
 Following directly from the definitions of joint and conditional entropy, this proof illustrates how the chain rule naturally arises from these definitions.
@@
 (19) appears naturally in the final step, the chain rule formula is sometimes alternatively used as the definition of conditional entropy; see, for example, reference [Nielsen2012ch11].
-For the second version of the chain rule, we generalize to $n$ random variables.
-By repeatedly applying the two-variable chain rule, the entropy of $n$ random variables can be expressed as the sum of conditional entropies by induction [Cover2005].
+For the second version of the chain rule, we generalize to $n$ random variables through induction by repeatedly applying the two-variable chain rule.
 Theorem 2 (Chain rule: $n$ random variables).
-$S(X_{1},X_{2},...,X_{n})=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1})$ (28)
+The entropy of $n$ random variables can be expressed as the sum of conditional entropies
+$S(X_{1},X_{2},\ldots,X_{n})=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1}).$ (27)
 Proof.
@@
 $=\vdots$ (31)
-$S(X_{1},X_{2},...X_{n})$ $=S(X_{1})+S(X_{2}\mid X_{1})+...+S(X_{n}\mid X_{n-1},...X_{1})$ (33)
-$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1})\qed$ (34)
-Thus far, we have naturally treated each random variable on an equal level in accordance with the entropy being a symmetric function of the random variables.
-However, it is sometimes useful to consider one (or more) random variables hierarchically.
-That is, we wish to consider the entropy between configurations and the entropy within those configurations.
-To do this, we first define the entropy within a given configuration $k$:
+$S(X_{1},X_{2},\ldots X_{n})$ $=S(X_{1})+S(X_{2}\mid X_{1})+\ldots+S(X_{n}\mid X_{n-1},\ldots,X_{1})$ (32)
+$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1})\qed$ (33)
+Thus far, we have treated each random variable on the same hierarchy, and it is natural to do so as the Shannon entropy is symmetric under permutation of the arguments.
+However, it is sometimes useful to consider one or more random variables hierarchically.
+That is, to treat some random variables as superior to others, and examine the entropy that remains when their outcomes are fixed.
+To do this, let $K$ be a random variable with probability mass function $p(k)$, where $\sum_{k\in\mathcal{K}}p(k)=1$.
 Definition 4.
-$S_{k}$ $=S(X_{1},X_{2},...X_{n}\mid K=k)$ (35)
-$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1},K=k)$ (36)
-Let $K$ be a random variable with probability distribution $p(K)$, where $\sum_{k}p_{k}=1$. We are now ready to prove our third version of the chain rule for the case of multiscale entropy.
+The entropy within a given configuration $k$, called the intra-configurational entropy is defined as
+$S_{k}$ $=S(X_{1},X_{2},\ldots,X_{n}\mid K=k)$ (34)
+$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1},K=k).$ (35)
+In this way, we can treat $K$ as superior to the other random variables $X$, and all the fully-constrained $X$-configurations are grouped by the outcomes $k$.
+However, note that the groups need not form a partition of the fully-constrained configurations, contrasting with Section 3, since different $k$ values may assign nonzero probability to the same $X$-configuration.
+We are now ready to prove our third version of the chain rule for the case of multiscale entropy.
 Theorem 3 (Recursive property).
-$S(K,X_{1},X_{2},...,X_{n})=-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}$ (37)
+The entropy of $n+1$ random variables may be expressed recursively as
+$S(K,X_{1},X_{2},\ldots,X_{n})=-\sum_{k\in\mathcal{K}}p(k)\ln p(k)+\sum_{k\in\mathcal{K}}p(k)S_{k}.$ (36)
+For convenience, we may define $p_{k}=p(k)$ and write the sum over $k\in\mathcal{K}$ simply as $\sum_{k}$:
+$\boxed{S(K,X_{1},X_{2},\ldots,X_{n})=-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}.}$ (37)
 Proof.
-First we choose to write the chain rule in the following form using the definition of conditional entropy from equation 19.
-$S(X_{1},X_{2},...,X_{n})$ $=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1})$ (38)
-$S(K,X_{1},X_{2},...,X_{n})$ $=S(K)+\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1},K)$ (39)
-$=S(K)+\sum_{k\in\mathcal{K}}p(k)\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1},K=k)$ (40)
-$=-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}\qed$ (41)
-Note that in this form, there is a conspicuous absence of any reference to any of the $X_{i}$ on the right-hand side as they are all subsumed into $S_{k}$.
+We begin with the chain rule:
+$S(X_{1},X_{2},\ldots,X_{n})=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1}).$ (38)
+Then, we introduce $K$ and use the chain rule to separate the entropy of $K$ from the remaining variables.
+We then apply the definitions of conditional entropy (Eq.
+(18)) and intra-configurational entropy $S_{k}$ (Eq.
+(35)):
+$S(K,X_{1},X_{2},\ldots,X_{n})$ $=S(K)+\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1},K)$ (39)
+$=S(K)+\sum_{k\in\mathcal{K}}p(k)\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1},K=k)$ (40)
+$=-\sum_{k\in\mathcal{K}}p(k)\ln p(k)+\sum_{k\in\mathcal{K}}p(k)S_{k}.\qed$ (41)
+Note that in this form, there is a conspicuous absence of any reference to any of the $X_{i}$ on the right-hand side of the last equality, as they are all subsumed into $S_{k}$.
 Hence, the recursive property can be regarded as the hierarchical form of the chain rule.
-The derivations presented in this section establish the mathematical foundation of the recursive property and show that the recursive property is not only consistent with the definition of entropy, but also essential.
-In fact, the recursive property is often taken as an axiom to single out the Shannon entropy as the natural choice of equations [khinchin1953concept, faddeev1956concept, Aczel1974] for entropy.
-The recursive property, as a hierachical form of the chain rule, reveal that the total entropy can be decomposed into a configuration component $\sum_{k}p_{k}\ln p_{k}$ and weighted sum of the intra-configurational entropies $\sum_{k}p_{k}S_{k}$.
-The recursive property represents a generalization of the standard entropy definition that reduces to the familiar form when configurations are chosen to be quantum pure states.
-That is, if $k$ specifies all degrees of freedom any other information ($X_{i}$) willbe redundant and the intra-configurational entropy is zero for all $k$, $\forall k\in\mathcal{K},\;S_{k}=0$.
-This flexibility in choosing the scale of description provides a powerful multiscale framework known as Zentropy for addressing systems on multiple scales where a natural partitioning of configurations exists, enabling efficient coarse-graining of degrees of freedom while maintaining thermodynamic consistency.
-5 Derivation of the partition function
-Definition 5.
-The Helmholtz energy is defined as
-F=ETSF=E-TS (42)
-where the energy of the thermodynamic system is the expectation value of the fully specified configurations
-$E=\langle E_{x}\rangle=\sum_{x}p_{x}E_{x}$ (43)
-For the case where partially specified configurations form a partition, as we did in the illustrative example, by the law of total expectation, we know that $\langle E_{x}\rangle=\langle E_{k}\rangle$, and the energy of the system can also be expressed as the expectation value over the partially specified configurations.
-$E=\sum_{k}p_{k}E_{k}$ (44)
-As each configuration is now treated as an ensemble, we would expect $F_{k}$ to take the place of $E^{k}$ in the standard approach.
-Indeed, this is what we find.
-Substituting equations Eq.
-(44) and Eq.
-(37) into the definition of Helmholtz energy gives
-$F=\sum_{k}p_{k}E_{k}+T\left(\sum_{k}p_{k}\ln p_{k}-\sum_{k}p_{k}S_{k}\right)$ (45)
-Then by the definition of Helmholtz energy Def. 5, $F_{k}=E_{k}-TS_{k}$, and Eq. (45) simplies to
-$F=\sum_{k}p_{k}F_{k}+T\sum_{k}p_{k}\ln p_{k}$ (46)
-With the thermodynamic system energy and Helmholtz energies now given in terms of $E_{k}$ and $S_{k}$ we can now derive the partition function in terms of $E_{k}$ and $S_{k}$.
-Following the standard derivation of the partition function, we maximize entropy $S=-\sum_{k}p_{k}\ln p_{k}+\sum p_{k}S_{k}$ under the constraints that the sum of the probabilities is equal to one
-$\sum_{k}p_{k}=1$ (47)
-and the internal energy is a constant
-$U=\langle E\rangle=\sum_{k}p_{k}E_{k}$ (48)
-The Lagrange function is
-$\mathcal{L}=\left(-\sum_{k}p_{k}\ln p_{k}+\sum p_{k}S_{k}\right)+\alpha\left(1-\sum_{k}p_{k}\right)+\beta\left(U-\sum_{k}p_{k}E_{k}\right)$ (49)
-To extremize entropy we set the variation of the Lagrange function to zero
+The derivations in this section establish the mathematical foundation of the zentropy and demonstrate the consistency of the recursive property with the definition of Shannon entropy.
+This property is fundamental and often used as an axiom to uniquely determine the Shannon entropy as the most natural measure of uncertainty [khinchin1953concept, faddeev1956concept, Aczel1974].
+The recursive property, as a hierarchical form of the chain rule, reveals that the total entropy can be decomposed into configurational and intra-configurational terms:
+$S=S_{\text{conf}}+\langle S_{k}\rangle,\\$ (42)
+where the configurational entropy is the entropy on the preferred level of hierarchy defined by $K$ and the intra-configurational term is the expectation value of the intra-configurational entropies:
+$S_{\text{conf}}$ $=-\sum_{k\in\mathcal{K}}p(k)\ln p(k)$ (43)
+$\langle S_{k}\rangle$ $=\sum_{k\in\mathcal{K}}p(k)S_{k}.$ (44)
+The recursive property provides a hierarchical decomposition of the standard entropy definition that reduces to the familiar form when configurations are fully-constrained with no microscopic degrees of freedom.
+That is, if $k$ specifies all degrees of freedom, any other information ($X_{i}$) will be redundant, and the intra-configurational entropy is zero for all $k$ (that is, $\forall k\in\mathcal{K},\;S_{k}=0$).
+This flexibility in choosing the scale of description provides a powerful multiscale framework known as zentropy for addressing systems on multiple scales, enabling efficient coarse-graining of degrees of freedom while maintaining thermodynamic consistency.
+5 Helmholtz energy and the Partition Function
+Since the Helmholtz energy and the partition function are defined downstream from entropy, we have to take into account intra-configurational entropy when evaluating these quantities for partially-constrained configurations.
+Since the partially-constrained configurations contain both energy and entropy, we can think of the partially-constrained configurations as sub-ensembles.
+In fact, in this section, we show that the Helmholtz energy and partition function are evaluated in terms of an intrinsic Helmholtz energy $F_{k}$ for the partially-constrained case instead of energy $E_{x}$ as in the fully-constrained case.
+From thermodynamics, the Helmholtz energy is defined as
+F=UTS,F=U-TS, (45)
+where the internal energy $U$ of the thermodynamic system is the expectation value of the energy of the fully-constrained configuration:
+$U=\langle E_{x}\rangle=\sum_{x}p_{x}E_{x}.$ (46)
+For the case where the partially-constrained configurations form a partition, as we did in the illustrative example in Section 3, the law of total expectation applies so $\langle E_{x}\rangle=\langle E_{k}\rangle$, and the energy of the system can also be expressed as the expectation value over the partially-constrained configurations:
+$U=\sum_{k}p_{k}E_{k}.$ (47)
+Naively, we might expect that $F=\langle F\rangle=\sum_{k}p_{k}F_{k}$. However, substituting Eq. (47) and Eq. (37) into the definition of Helmholtz energy (Eq. (45)) gives
+$F=\sum_{k}p_{k}E_{k}+T\left(\sum_{k}p_{k}\ln p_{k}-\sum_{k}p_{k}S_{k}\right).$ (48)
+For convenience, we define the quantity $F_{k}=E_{k}-TS_{k}$, and Eq. (48) simplifies to
+$\boxed{F=\sum_{k}p_{k}F_{k}+T\sum_{k}p_{k}\ln p_{k}.}$ (49)
+Indeed, for partially-constrained configurations that form a partition, the Helmholtz energy $F_{k}$ replaces the energy $E_{x}$ of fully-constrained configurations.
+Next, we derive the partition function by maximizing entropy in its recursive form (Eq. (37)) under the constraints that the sum of the probabilities is equal to one,
+$\sum_{k}p_{k}=1,$ (50)
+and that the expectation value of the energy is a constant,
+$U=\sum_{k}p_{k}E_{k}.$ (51)
+Following the method of Lagrange multipliers in the calculus of variations, the Lagrange function is
+$\mathcal{L}=\left(-\sum_{k}p_{k}\ln p_{k}+\sum p_{k}S_{k}\right)+\alpha\left(1-\sum_{k}p_{k}\right)+\beta\left(U-\sum_{k}p_{k}E_{k}\right).$ (52)
+To extremize entropy, we set the variation of the Lagrange function to zero and simplify:
 $0$ $=\delta\mathcal{L}$ (53)
 $=\delta\left(-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}\right)+\delta\left(\alpha-\sum_{k}\alpha p_{k}\right)+\delta\left(\beta U-\sum_{k}\beta p_{k}E_{k}\right)$ (54)
-$=\sum_{k}\left[\delta\left(-p_{k}\ln p_{k}+p_{k}S_{k}\right)-\delta\left(\alpha p_{k}\right)-\delta\left(\beta E_{i}p_{i}\right)\right]$ (52)
-$=\sum_{k}\left[\frac{\partial}{\partial p_{k}}\left(-p_{k}\ln p_{k}\right)\delta p_{k}+\frac{\partial}{\partial p_{k}}\left(p_{k}S_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\alpha p_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\beta E_{i}p_{i}\right)\delta p_{k}\right]$ (53)
-$=\sum_{k}\left[-\ln p_{k}-1+S_{k}-\alpha-\beta E_{i}\right]\delta p_{k}$ (54)
-relying on the fact that entropy is symmetric with respect to $p_{k}$, for any variation
-$0=-\ln p_{k}-1+S_{k}-\alpha-\beta E_{k}$ (55)
+$=\sum_{k}\left[\delta\left(-p_{k}\ln p_{k}+p_{k}S_{k}\right)-\delta\left(\alpha p_{k}\right)-\delta\left(\beta E_{k}p_{k}\right)\right]$ (55)
+$=\sum_{k}\left[\frac{\partial}{\partial p_{k}}\left(-p_{k}\ln p_{k}\right)\delta p_{k}+\frac{\partial}{\partial p_{k}}\left(p_{k}S_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\alpha p_{k}\right)\delta p_{k}-\frac{\partial}{\partial p_{k}}\left(\beta E_{k}p_{k}\right)\delta p_{k}\right]$ (56)
+$=\sum_{k}\left[-\ln p_{k}-1+S_{k}-\alpha-\beta E_{k}\right]\delta p_{k}.$ (57)
+Since the variations are independent, the coefficient of each $\delta p_{k}$ must vanish. Thus, for every $k$,
+$0=-\ln p_{k}-1+S_{k}-\alpha-\beta E_{k}.$ (58)
 Isolating $p_{k}$ yields
-$p_{k}=e^{\left(-1-\alpha-\beta E_{k}+S_{k}\right)}$ (56)
-applying our probability constraint Eq.(47)
+$p_{k}=e^{\left(-1-\alpha-\beta E_{k}+S_{k}\right)}.$ (59)
+Applying the probability constraint (Eq. (50)),
 $1$ $=\sum_{k}p_{k}$ (60)
@@
 $=e^{-(1+\alpha)}\sum_{k}e^{\left(-\beta E_{k}+S_{k}\right)}$ (63)
-$\frac{1}{e^{-(1+\alpha)}}=\sum_{k}e^{\left(-\beta E_{k}+S_{k}\right)}$ (61)
+$\frac{1}{e^{-(1+\alpha)}}=\sum_{k}e^{\left(-\beta E_{k}+S_{k}\right)}.$ (64)
 The partition function is then naturally defined as
-$Z=\frac{1}{e^{-(1+\alpha)}}$ (62)
+$Z=\frac{1}{e^{-(1+\alpha)}}.$ (65)
 Substituting into Eq. (64) gives the relation between the partition function and $\beta$:
 $Z=\sum_{k}e^{\left(-\beta E_{k}+S_{k}\right)}.$ (66)
-While we have derived the partition function in terms of $E_{k}$ and $S_{k}$, in order to write it in terms of $F_{k}$, we assume here that $\beta$ is still inverse temperature then prove it subsequently.
-$Z$ $=\sum_{k}e^{\left(-\frac{1}{T}E_{k}+S_{k}\right)}$ (64)
-$=\sum_{k}e^{\left(-\frac{1}{T}(E_{k}-TS_{k})\right)}$ (65)
-$=\sum_{k}e^{\left(-\frac{F_{k}}{T}\right)}=\sum_{k}e^{\left(-\beta F_{k}\right)}$ (66)
-To confirm that the identity of $\beta$ is inverse temperature and is unaffected by the inclusion of intra-configurational entropy, we write $p_{k}$ in terms of the partition function in order to write entropy in terms of the partition function and finally apply the definition of temperature.
+Assuming, that $\beta$ is inverse temperature, we can write the partition function in terms of $F_{k}$,
+$\boxed{Z=\sum_{k}e^{\left(-\frac{F_{k}}{T}\right)}=\sum_{k}e^{\left(-\beta F_{k}\right)}.}$ (67)
+To confirm that the identity of $\beta$ is indeed inverse temperature and is unaffected by the inclusion of intra-configurational entropy, we write $p_{k}$ and entropy $S$ in terms of the partition function and then apply the definition of temperature.
 Substituting Eq.
@@
 (59),
-$p_{k}=\frac{1}{Z}e^{\left(-\beta E_{k}+S_{k}\right)}$ (68)
-Substituting into the definition of the logarithmic portion of entropy
+$p_{k}=\frac{1}{Z}e^{\left(-\beta E_{k}+S_{k}\right)}.$ (68)
+Next, we rewrite the recursive property (Eq. (37)) using Eq. (68) to evaluate $\ln p_{k}$ (leaving the prefactor and other instances of $p_{k}$ unchanged:
 $S$ $=-\sum_{k}p_{k}\ln\left[\frac{1}{Z}e^{\left(-\beta E_{k}+S_{k}\right)}\right]+\sum_{k}p_{k}S_{k}$ (69)
@@
 $=\beta\sum_{k}p_{k}E_{k}+\ln{Z}\sum_{k}p_{k}$ (74)
-$=\beta E+\ln{Z}$ (75)
-Differentiating with respect to energy gives
-$\frac{\partial S}{\partial E}_{V,N}=\beta$ (77)
-and by the definition of temperature $T=\frac{\partial E}{\partial S}_{V,N}$,
-$\beta=\frac{1}{T}$ (78)
+$=\beta U+\ln{Z}.$ (75)
+Notice that the $S_{k}$ terms cancel out in Eq. (71). Differentiating with respect to the internal energy gives
+$\left(\frac{\partial S}{\partial U}\right)_{V,N}=\beta,$ (76)
+and by the definition of temperature $T=\left(\frac{\partial U}{\partial S}\right)_{V,N}$,
+$\beta=\frac{1}{T},$ (77)
 as expected.
-Note that the definition of the partition function eq.
-(62) and the identity of $\beta$ are the same as in the standard thermodynamic approach, the derived relation between the partition function and $\beta$ is not.
-The difference being that $E_{x}$ is replaced by $F_{k}$, as expected.
+Note that while the definition of the partition function (Eq.
+(65)) and the identity of $\beta$ (Eq.
+(77)) are the same as in the standard thermodynamic approach, the derived relation between the partition function and $\beta$ (Eq.
+(67)) is not.
+The difference being that energy is replaced by Helmholtz energy, as expected.
 6 Usefulness of Zentropy
-Why do we need recursive entropy when the definition is sufficient?
-Why group pure states?
-The answer is that recursive entropy is a generalization of the definition that provides a natural and efficiently computable approach to commonly encountered problems.
-Namely, problems where states of the system may be partitioned into basins.
-This partitioning into basins explains emergent pheonomena.
-Thus a framework, such as zentropy, that incorporates this partitioning of states, naturally predicts emergent phenomena.
-The recursive property of entropy (Eq.
-1) is a generalization of the definition.
-If the set of states are partitioned so finely that each $k$ already specifies a pure state (the number of $k$ is maximal), then $S_{k}$ is equal to zero for all $k$, only the first term $k_{B}\sum_{k}p_{k}\ln p_{k}$ remains, $k=x$, and we recover the definition of entropy (Eq.
-2).
-Keeping the intra-configurational term therefore just allows us to choose configurations that contain additional unresolved mixed-up-ness.
-Recursive entropy is thus a generalization and reduces to the definition when no hidden uncertainty remains in each configuration ($S_{k}=0)$.
-Conversely, at the opposite extreme, we could choose our set of states to be as coarsely partitioned as possible so that there is only one configuration (The number of $k$ is minimal), in which case, the configurational entropy term is zero $k_{B}\sum_{k}p_{k}\ln p_{k}=0$ and Eq.
-1 becomes trivial $S=\sum_{k}p_{k}S_{k}=S_{k}$.
-However, it is often best to choose a partitioning that is somewhere inbetween.
-All this is to say that by including the intra-configurational entropy term $\sum_{k}p_{k}S_{k}$—using the recursive property of entropy— allows us to choose a definition of configuration that is most useful.
-Or in other words, to define the configurations of the system at the most useful scale.
-A useful way to group pure states is in the same way as they are partitioned in nature, that is, to partition pure states in accordance with the basins of the energy landscape.
-This way, it is natural to consider the system’s behavior given a specifed basin/configuration.
-This can be used to simplify calculations by applying approximations hierarchically, as the pure states in each basin often have some shared properties.
+Why do we need recursive entropy (Eq.
+(2)) when we have the standard definition (Eq.
+(1))?
+Why group microstates at all?
+The answer is that recursive entropy is a generalization of the standard definition that provides a natural and computationally efficient approach for commonly encountered problems.
+Namely, problems where microstates of the system may be partitioned into basins.
+This partitioning into basins explains emergent phenomena.
+Thus, a framework, such as zentropy, that incorporates a hierarchical partitioning of microstates, naturally predicts emergent phenomena.
+The recursive property of entropy reduces to the standard definition in the limit of fully-constrained configurations.
+If the set of microstates is partitioned so finely that each $k$ already specifies a fully-constrained configuration (the number of $k$ is maximal), then $S_{k}$ is equal to zero for all $k$, only the first term $k_{B}\sum_{k}p_{k}\ln p_{k}$ remains.
+Identifying $k=x$, we recover the usual definition of entropy.
+Conversely, at the opposite extreme, we could choose our set of microstates to be as coarsely partitioned as possible, so that there is only one configuration (the number of $k$ is minimal), in which case, the configurational entropy term vanishes ($k_{B}\sum_{k}p_{k}\ln p_{k}=0$) and Eq.
+(2) becomes trivial $S=\sum_{k}p_{k}S_{k}=S_{k}$.
+In practice, however, the most useful description typically lies between these two extremes.
+All this is to say that including the intra-configurational entropy term $\sum_{k}p_{k}S_{k}$ allows us to choose a definition of configuration that is best suited to the problem at hand.
+Or, in other words, to define the configurations of the system at the most useful scale.
+We emphasize that the recursive property is not a statement about the intrinsic nature of a system; rather, its usefulness lies in the flexibility it gives us in describing the system in terms of partially-constrained configurations.
+A useful way to group microstates is in the same way as they are partitioned in nature, that is, to partition microstates in accordance with the basins of the energy landscape.
+This way, it is natural to consider the system’s behavior given a specified basin/configuration.
+This can be used to simplify calculations by applying approximations hierarchically, as the microstates in each basin often have shared properties.
 For example, consider a magnetic crystal, where each configuration is a given magnetic ordering.
-The remaining uncertainty is then in the phonon and electron states given that magnetic ordering.
-$S_{k}=S_{k,\text{el}}+S_{k,\text{vib}}$ (79)
+The remaining uncertainty is then in the phonon and electron states given that magnetic ordering:
+$S_{k}=S_{k,\text{el}}+S_{k,\text{vib}},$ (78)
 where $S_{k}$ is the entropy of a given configuration/magnetic ordering, $S_{\text{el}}$ is entropy due to electron excitations, and $S_{\text{vib}}$ due to lattice vibrations.
@@
 This assumption facilitates the calculations.
-Otherwise, they are subadditive, and an additional relation between $S_{\text{el}}$ and $S_{\text{vib}}$ is required to determine the $S_{k}$ exactly.
-However, such a calculation is difficult and often unnecessary.
-Here we clarify some terminology.
-There are two different uses of the term ”coarse-graining”.
-To clarify, one meaning refers to the partitioning or, more generally, grouping of states, and the other refers to the incomplete sampling of configurations.
-While the term ”coarse-graing” is often used for both, The two meanings are very different.
-Notably, incomplete sampling is approximate, whereas grouping of states and calculating entropy with the recursive property is exact.
+Otherwise, they are subadditive, and an additional relation between $S_{k,\text{el}}$ and $S_{k,\text{vib}}$ is required to determine $S_{k}$ exactly.
+However, such a calculation is difficult and often unnecessary [Shang2010b, Wang2004].
+Here, we clarify some terminology.
+There are two different uses of the term “coarse-graining”.
+The first, which we have already dealt with, refers to the partitioning or, more generally, grouping of microstates, while the other refers to the incomplete sampling of configurations.
+Although the term “coarse-graining” may be used for both [Bengtsson2017, Kmiecik2016], the two meanings are fundamentally different.
+Notably, incomplete sampling is approximate, whereas grouping of microstates and calculating entropy with the recursive property is exact.
 We now turn our attention to the incomplete sampling of configurations.
-While approximate, it is often not feasible to calculate all possible configurations of a system as required by the partition function.
+Although approximate, it is often infeasible to calculate all possible configurations of a system, as required by the partition function.
 Instead, a representative sample of configurations may be used.
-For example, this may be achieved by calculating all the configurations of a finite cell rather than the infinite lattice, or using Monte Carlo or other methods to sample only configurations of sufficient probability.
-By grouping pure states, we can more efficiently sample or approximate the pure states.
-For example, by partitioning pure states by spin configuration, we can apply the harmonic approximation to each configuration to gather the phonon density of states.
-Including both configurational and intra-configurational terms in the equation for entropy, allows defining configurations on a scale between the macroscopic thermodynamic state and quantum pure states.
-As mentioned earlier, for partitions, when the number of configurations $k$ is maximal, the configurations are defined on the scale of quantum pure states, and when the number of configurations is minimal, the configuration is the macroscopic thermodynamic state.
-By choosing the partition such that the number of configurations is between the two extremes, we build a bridge between scales.
-This bridge can be useful in the description of emergent phenomena.
-While there are rules that govern the quantum pure states (quantum mechanics), one could form new rules that govern the configurations formed by the grouping.
-Thus, a rule set for the quantum pure states, and an emergent rule set for the configurations govern the same system.
-This approach works particularly well for configurations wherein the quantum pure states (or more gennerally subconfigurations) are strongly lumpable [Rosas2024].
-That is, the quantum pure states of one group have a defining property that is clearly distinct from quantum pure states of another configuration.
+For example, this may be achieved by calculating all the configurations of a finite cell rather than the infinite lattice, or by employing Monte Carlo or related methods to sample only configurations with sufficiently high probability.
+Including both configurational and intra-configurational terms in the equation for entropy allows defining configurations on a scale between the macroscopic thermodynamic state and fully-constrained microscopic pure states.
+Thus, we build a bridge between scales, as illustrated in Figure 1b, which can be useful in the description of emergent phenomena.
+While there are rules that govern the fully-constrained configurations (quantum mechanics), one can define new, emergent rules that govern the grouped configurations.
+In this way, a system can be described both by the rule set for fully-constrained configurations and by an emergent rule set for the grouped configurations.
+This approach works particularly well for configurations wherein the subconfigurations are strongly lumpable [Rosas2024].
+That is, the subconfigurations of one configuration have a defining property that is clearly distinct from the subconfigurations of another configuration.
+Ultimately, the utility of zentropy lies in the ability to describe systems accurately despite the practical necessity of incomplete sampling.
+By providing a mathematical structure that adapts to the natural basins of the energy landscape, the framework renders microscopic complexity tractable on macroscopic scales.
+This flexibility streamlines computation and illuminates the causal links between fundamental interactions and emergent macroscopic behavior.
+7 Conclusion
 In summary, the recursive property of entropy provides a flexible and exact framework for partitioning thermodynamic systems at physically meaningful scales.
-By grouping pure states according to natural basins in the energy landscape, we can apply hierarchical approximations that exploit shared properties within each configuration, simplifying otherwise intractable calculations.
-The multiscale nature of the approach naturally bridges microscopic and macroscopic descriptions.
-This bridging capability makes zentropy particularly powerful for describing emergent phenomena, where different rule sets govern behavior at different scales, provided the configurations are sufficiently distinct or ”strongly lumpable.” The result is a computationally practical approach that maintains rigorous statistical-mechanical foundations while enabling quantitative predictions across scales.
+By grouping microstates according to natural basins in the energy landscape, we can apply hierarchical approximations within each configuration, exploiting the shared properties of the microstates and thereby simplifying otherwise intractable calculations.
+The multiscale nature of the approach naturally bridges microscopic and macroscopic descriptions, which makes zentropy particularly powerful for describing emergent phenomena, where different rule sets govern behavior at different scales, provided the subconfigurations are sufficiently distinct or “strongly lumpable.” The result is a computationally practical approach that maintains rigorous statistical-mechanical foundations while enabling quantitative predictions across scales.
 Here, an analogy can be drawn between zentropy and Kohn–Sham DFT, which provides a practical approach to quantum mechanics by replacing an intractable interacting many-electron problem with a non-interacting system in an effective potential that preserves the exact density.
-7 Conclusion
-We have established a clear statistical-mechanical foundation for the zentropy approach by proving the recursive property of entropy as a hierarchical form of the chain rule and showing its equivalence to the standard definition when configurations reduce to pure states.
+We have established a clear statistical-mechanical foundation for the zentropy approach by proving the recursive property of entropy as a hierarchical form of the chain rule and showing its equivalence to the standard definition when configurations are fully constrained.
 Building on this foundation, we derived a grouped form of the Helmholtz energy and obtained a partition function expressed in terms of configuration Helmholtz energies.
-This zentropy approach separates configurational entropy from intra-configurational contributions and provides a principled, exact route to multiscale modeling, by aligning configurations with natural basins in the energy landscape
-Practically, these results position zentropy as a computationally efficient bridge between microscopic states and macroscopic behavior, clarifying when and how emergent phenomena can be captured by a reduced description.
-We anticipate that adopting the recursive property in thermodynamics and condensed-matter calculations will streamline predictive workflows, particularly in DFT and when configurations are strongly lumpable.
+This zentropy approach separates configurational entropy from intra-configurational contributions and provides a principled, exact route to multiscale modeling, by aligning configurations with natural basins in the energy landscape.
+Practically, these results position zentropy as a computationally efficient bridge between microscopic and macroscopic behavior, clarifying when and how emergent phenomena can be captured by a reduced description.
+We anticipate that adopting the recursive property in thermodynamics and condensed-matter calculations will streamline predictive workflows, particularly in DFT and when subconfigurations are strongly lumpable.
 8 Acknowledgments
-The present work is funded by the Department of Energy (DOE) through Grant No. DE-SC0023185
+The present work is funded by the Department of Energy (DOE) through Grant No. DE-SC0023185.
+CRediT author statement
+Luke Allen Myers: Conceptualization, Methodology, Writing - Original Draft, Writing - Review & Editing, Visualization
+Nigel Lee En Hew: Conceptualization, Writing - Review & Editing, Supervision
+Shun-Li Shang: Writing - Review & Editing, Supervision, Funding acquisition
+Zi-Kui Liu: Writing - Review & Editing, Supervision, Funding acquisition